# Calibration Attack: A Framework For Adversarial Attacks Targeting Calibration

### v1

@@ Abstract
We introduce a new framework of adversarial attacks, named calibration attacks, in which the attacks are generated and organized to trap victim models to be miscalibrated without altering their original accuracy, hence seriously endangering the trustworthiness of the models and any decision-making based on their confidence scores. Specifically, we identify four novel forms of calibration attacks: underconfidence attacks, overconfidence attacks, maximum miscalibration attacks, and random confidence attacks, in both the black-box and white-box setups. We then test these new attacks on typical victim models with comprehensive datasets, demonstrating that even with a relatively low number of queries, the attacks can create significant calibration mistakes. We further provide detailed analyses to understand different aspects of calibration attacks. Building on that, we investigate the effectiveness of widely used adversarial defences and calibration methods against these types of attacks, which then inspires us to devise two novel defences against such calibration attacks.
@@ 1 Introduction
Figure 1: Calibration diagrams of a normal ResNet-50 classifier before and after three forms of calibration attacks. Red bars show the average accuracy on the test data binned by confidence scores and blue bars represent the actual average confidence in each bin. The yellow line depicts perfect calibration. Despite the accuracy being unchanged, the miscalibration is severe after the attacks.
Adversarial attacks (Ren et al., 2020) have proven to be a critical tool to reveal the susceptibility of deep neural networks  (Ibitoye et al., 2019; Zimmermann et al., 2022; Xiao et al., 2023), where, in a typical setup, adversarial examples are generated by introducing imperceptible perturbations on an original input to cause model misclassification. The existing attacks have mainly focused on trapping victim models to make incorrect predictions.
In this paper, we highlight a serious set of threats, named Calibration Attack, which focus on attacking the calibration of victim models by deceiving the models to be over or under-confident without modifying their prediction accuracy, hence endangering any decision-making based on a classifier’s confidence scores. We briefly illustrate this in Figure 1.
Calibration attacks are insidious and arguably harder to detect than standard attacks, which we empirically demonstrate in Table 2, when using popular adversarial attack detection techniques. The harm presented from these attacks is that on the surface the model appears to still be making mostly correct decisions, but the level of miscalibration could make the model’s decisions unusable for downstream decision making. Consider the case of medical diagnosis, where after a set of X-ray scans are calibration-attacked, it can appear as though a diagnosis model is performing as usual, but the confidence levels on ambiguous cases that need further review by an expert will be altered, and hence compromise the affected decision-making leading to potential complications.
In our study, we propose four forms of calibration attacks: underconfidence attack, overconfidence attack, maximum miscalibration attack, and random confidence attack. We test these newly identified attacks primarily in the black-box setup due to it being a more realistic and applicable threat scenario, but we also test white-box variations. We use benchmark image classification datasets to evaluate the models. We also analyze the adequecy of existing adversarial defenses and calibration techniques against maintaining good calibration post-attack, as well as formulate and study two novel forms of defence, Calibration Attack Adversarial Training (CAAT) and Compression Scaling (CS), helping to bring additional insights for how to best resist these calibration attacks.
Following related attacks on credible uncertainty estimates (Galil & El-Yaniv, 2021), we build on previous works by providing a comprehensive investigation on numerous forms of calibration attacks, especially on the threat of inducing overconfidence and their effectiveness on defenses. In summary, our main contributions are as follows. 1) We introduce and investigate a new form of adversarial attack, the calibration attack, and advocate that particular caution should be seriously taken for such attacks. 2) We construct four specific forms of calibration attacks and show their danger to typical deep learning models. Detailed analyses are provided on the effectiveness of different attacks and the vulnerability of different architectures of victim models. 3) We show the severity of calibration attacks on many models using current calibration and defence methods, and demonstrate the compromise often presented between robustness and calibration, but offer additional insights for future defences based on our two novel methods.
@@ 2 Related Work
Calibration of Machine Learning Models. Calibration methods are generally divided into two types. Post calibration can be applied directly to the predictions of fully trained models at test time, which include classical approaches like temperature scaling (Guo et al., 2017), Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002), and histogram binning (Zadrozny & Elkan, 2001). Training-based approaches, however, typically add a bias during training to help a model to be better calibrated (Zhang et al., 2018; Thulasidasan et al., 2019; Kumar et al., 2018; Tomani & Buettner, 2021). In this work we investigate how a diverse range of calibration methods cope against calibration attacks and study the limitations that require them to be overhauled to deal with attacks, including the vulnerability of models on convolutional (Guo et al., 2017; Minderer et al., 2021), and Transformer architecture ViT (Dosovitskiy et al., 2021).
Adversarial Attacks and Training. We provide a more detailed overview of adversarial attack methods in Appendix A. Black-box attacks  (Carlini & Wagner, 2017) have less information about victim models. Many methods of this type are based on some form of gradient estimation through querying the model and finding the finite differences (Bhagoji et al., 2018). In contrast, white-box attacks often have access to the full details of a victim model such as model architectures and gradients. For example, a typical example of this type, among many others, is the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015).
In terms of adversarial training, many approaches have also been introduced (Stutz et al., 2020; Chen et al., 2022; Qin et al., 2021; Patel et al., 2021; Dhillon et al., 2018). For example, Stutz et al. (2020) improve the robustness of the models against previously unseen attacks by rejecting low-confidence adversarial examples through confidence-thresholding; Chen et al. (2022) design a post processing adversarial defence method against score-based query attacks. In addition, previous works on defense methods that certify the adversarial robustness have certified confidence scores of images. In Kumar et al. (2020), they leverage information about the distribution of confidence scores around an input image to generate certified radii for the prediction confidence of the smoothed classifier. Building on this, in Emde et al. (2023), they propose certifying the calibration by generating the worst-case bounds of calibration error under potential adversarial perturbations and discuss the importance studying attacks targeting calibration metrics.
@@
Attacking Uncertainty Estimates. A few previous approaches examine attacks on confidence. Galil & El-Yaniv (2021) first identified attacks on credible uncertainty estimates as an issue. Their attacks formulated under both black box and white box settings, harm the potential of using uncertainty estimation on a model’s predictions by pushing correct samples closer to the decision boundary and incorrect ones further away using confidence without causing incorrect predictions. Zeng et al. (2023) introduce a data poisoning attack designed to alter the training process of models such that a high-confidence region forms around an out of domain sample, harming the adversary-resistant uncertainty estimates on a set of targeted out-domain samples while leaving the original labels. Nevertheless, none of the prior works comprehensively study the range of the four attacks we introduce and cover in the context of calibration rather than uncertainty estimation, nor do they systematically investigate how well victim models would remain well calibrated under such attacks, including under a range of different calibration methods and adversarial defense methods. Unlike previous works, we examine in detail the effects of attacks targeting overconfidence as well, and evaluate the detectability relative to standard attacks.
3 Calibration Attack
@@
3.1 Objective of Calibration Attack
Calibration attack aims to generate adversarial samples to maximize the miscalibration function $M(\tilde{\bm{x}},\bm{k})$ for a predicted class $k$ and an adversarial sample $\tilde{\bm{x}}$. As will be detailed in Section  3.2, we propose four specific forms of calibration attacks, in which the miscalibration function $M(\tilde{\bm{x}},\bm{k})$ has different implementations. The adversarial sample $\tilde{\bm{x}}$ is created by adding adversarial noise $\delta$ to an input vector $\bm{x}$: $\tilde{\bm{x}}=\bm{x}+\bm{\delta}$. To ensure that the adversarial example and the original do not differ a notable amount and following convention, a maximum $l_{p}$-ball norm difference bounded by an $\bm{\epsilon}$ is defined between them. For completeness of the paper, we include the details of adversarial constraint to be satisfied for certain $\bm{p}\in\{0,1,..,\infty\}$ norm:
@@ 3.1 Objective of Calibration Attack
$\begin{split}\|\tilde{\bm{x}}-\bm{x}\|_{\bm{p}}<\bm{\epsilon},\>\>s.t\>\>\hat{\bm{y}}\neq\bm{y},\end{split}$ (1)
where $\bm{\epsilon}$ controls the amount of perturbation allowed.
In general, our attacks are based on the most popular and easiest-to-measure view of calibration called classwise calibration that focuses on the predicted class (Kull et al., 2019). The equation below details the notion for all input datapoints $(\bm{x}_{n},\bm{y}_{n})\in\mathcal{D}=\{(\bm{x}_{n},\bm{y}_{n})\}^{N}_{n=1}$ in a dataset $\mathcal{D}$:
$\begin{split}\mathbb{P}(\bm{y}_{n}=\hat{\bm{y}}(\bm{x}_{n})\;|\;\hat{\bm{p}}_{k}(\bm{x}_{n})=\bm{p}_{\bm{k}})=\bm{p}_{\bm{k}},\end{split}$ (2)
@@
where $\bm{p}_{k}$ is the confidence score for the predicted class $\bm{k}$. Any mismatch between the left and right hand sides of the equation creates undesirable miscalibration.
3.2 Four Types of Calibration Attack
Within the aforementioned framework, we propose to construct four forms of calibration attack: underconfidence attack, overconfidence attack, maximum miscalibration attack, and random confidence attack. The first two types serve as the basis for the latter, while the maximum miscalibration and random confidence attacks are developed to create a higher and more problematic degree of miscalibration compared to the base attacks.
@@ 3.2 Four Types of Calibration Attack
Underconfidence and Overconfidence Attack. These two base attacks aim to solve the constrained optimization problem involving the miscalibration function $\bm{M}(\tilde{\bm{x}},\bm{k})$ for the predicted class $\bm{k}$. Specifically for the underconfidence attacks, the loss function is defined as:
$\begin{split}M(\tilde{\bm{x}},\bm{k})=\hat{\bm{p}}_{k}(\tilde{\bm{x}})-\max_{j\neq\bm{k}}\hat{\bm{p}}(\tilde{\bm{x}}).\end{split}$ (3)
The aim is to reduce the confidence scores on all predictions. For overconfidence attacks, the loss function is $M(\tilde{\bm{x}},\bm{k})=1-\hat{\bm{p}}_{k}(\tilde{\bm{x}})$, and the objective is updated so that adversarial examples are crafted to have the highest level of confidence. Both of these come with the added constraint that the examples are generated to try to keep predicted class label $\hat{\bm{y}}(\bm{x})=\hat{\bm{y}}(\tilde{\bm{x}})$, irrespective of true label $\bm{y}$, and with the goal of ensuring that the maximum amount of calibration error over the set of adversarial inputs $\tilde{\bm{\mathcal{X}}}$, $\;\mathop{\max}_{\tilde{\mathcal{X}}}(\mathbb{P}(\bm{y}_{n}=\hat{\bm{y}}(\tilde{\bm{x}}_{n})\;|\;\hat{\bm{p}}_{\bm{k}}(\tilde{\bm{x}}_{n})=\bm{p}_{\bm{k}})-\bm{p}_{\bm{k}})$ is produced.
The overall algorithm for the underconfidence and overconfidence attacks can be seen in Algorithm 1. The basis of our primary implementation of calibration attacks is the popular and highly effective Square Attack (SA) (Andriushchenko et al., 2020), a black-box score-based adversarial attack approach. It has attained state-of-the-art results in terms of query efficiency and success rate, even outperforming some white-box methods. Unlike the standard adversarial attack settings, we do not only run the attacks on correctly classified examples, but also the originally misclassified inputs, since they can also be further miscalibrated; attacking misclassified inputs in such a manner could be an additional avenue for adversaries to pursue to harm downstream decision-making. There are two basic forms of SA attacks, based on the $l_{\infty}$ or $l_{2}$ norm. We create a version of calibration attack for each norm. In addition to SA, we also create white-box variations of calibration attack using the effective Projected Gradient Descent (PGD) framework in the $l_{\infty}$ threat model. These attacks follow the same algorithm but have additional implementation details, which can be seen in Appendix B.
Maximum Miscalibration Attack. The main principle behind the maximum miscalibration attack is that to create the most amount of miscalibration in classifier for a given set of data, one needs to have all of the incorrectly classified samples have near 100% confidence, and all of the correctly classified samples have the minimum confidence scores.
1:  Input: Classifier $f$, input $\bm{x}$, original predicted class $\bm{k}$, $\bm{\epsilon}$, $I$ attack iterations, attack type $a$
2:  Output: Adversarial example $\tilde{\bm{x}}$
3:  $\tilde{\bm{x}}\leftarrow\bm{x}$
4:  if a=underconf then
5:     Loss function L returns the loss margin between predicted class and second highest class
6:  else if a=overconf then
7:     Loss function L returns softmax probability of predicted class
8:  end if
9:  $l^{*}\xleftarrow{}L(f(\tilde{x}),k)$
10:  for $i=1$ to $I$ do
11:     Perturb $\tilde{x}$ to find $\delta$ according to attack algorithm bounded by $l_{p}$ norm e.g. Square Attack
12:     $x_{new}\leftarrow$ Project $\tilde{x}+\delta$ to be within $\epsilon$ of the $l_{p}$ norm
13:     $l_{new}\leftarrow L(f(x_{new})),k)$
14:     if a=underconf and $l_{new}<l^{*}$ and $argmax(f(x_{new}))=k$ then
15:        $\tilde{x}\xleftarrow{}x_{new},l^{*}\xleftarrow{}l_{new}$
16:     else if a=overconf and $l_{new}>l^{*}$ then
17:        $\tilde{x}\xleftarrow{}x_{new}$, $l^{*}\xleftarrow{}l_{new}$
18:     end if
19:     if ($a$=underconf and $l^{*}<0.01$) or ($a$=overconf and $l^{*}>0.99$then
20:        break for loop
21:     end if
22:  end for
Algorithm 1 General Calibration Attack
Consider that, to achieve the highest amount of calibration error (100%) for a given set of data requires the following:
$\bm{p}_{k}=1,\hat{\bm{p}}_{k}(\bm{x}_{n})=1,\bm{y}_{n}\neq\hat{\bm{y}}(\bm{x}_{n})\forall\bm{x}_{n},$
Based on Eq. 2 , that will satisfy:
$\mathbb{P}(\bm{y}_{n}=\hat{\bm{y}}(\bm{x}_{n})\;|\;\hat{\bm{p}}_{k}(\bm{x}_{n})=\bm{p}_{k})=0.$
In other words, this occurs when a classifier only outputs 100% confidence scores, while its accuracy is 0.
Let a classifier have non-zero accuracy that is unchangeable. We cannot expect to reach the error of 100% since $\bm{p}_{k}=0$ cannot be the case (for the top predicted class) nor can $\mathbb{P}(\bm{y}_{n}=\hat{\bm{y}}(\bm{x}_{n}))=0$ be true for all $\bm{y}_{n}$. However, to achieve the highest calibration error on a set of data points in this scenario, one can first isolate the misclassified data points and if the classifier is made to output confidence scores of 100% on all of them, using calibration attack for example, it would create a total calibration error of 100% on this set of misclassified data points exactly as in the theoretical highest miscalibration scenario discussed previously. Next, with the remaining correctly classified points, which we know the classifier has an accuracy of 100% on, one can create the largest difference between the accuracy and average confidence by making the average confidence on this set as low as possible. Consider Eq. 2, since confidence scores can only range from $1/K$ to 1, the largest possible difference between the average confidence score and the accuracy of 100% is $1-1/K$. Again, if $\bm{p}_{k}=1/K$, and every $\hat{\bm{p}}_{k}(\bm{x}_{n})=1/K$, while $y_{n}=\hat{\bm{y}}(\bm{x}_{n})\forall\bm{x}_{n}$, then this makes the calibration error: $\mathbb{P}(\bm{y}_{n}=\hat{\bm{y}}(x_{n})\;|\;\hat{\bm{p}}_{k}(\bm{x}_{n})=\bm{p}_{k})-\bm{p}_{k}=1-1/K$. It is not possible to create a higher level of calibration error since if $\bm{p}_{k}>1/K$ on some number of the correctly classified samples, then $\mathbb{P}(y_{n}=\hat{\bm{y}}(\bm{x}_{n})\;|\;\hat{\bm{p}}_{k}(\bm{x}_{n})=\bm{p}_{k})$ will still be 1, while $\bm{p}_{k}>1/K$ will lead to less calibration error on that subset of samples. With errors on both mutually exclusive subsets of data maximized, the theoretical highest miscalibration will be created on the full data. Finally, to emulate this scenario, underconfidence attacks are conducted on all data points correctly classified by the model, and overconfidence attacks are conducted on all of the misclassified data yielding the maximum miscalibration attack.

### v2

# Calibration Attacks: A Comprehensive Study of Adversarial Attacks on Model Confidence

@@ Abstract
In this work, we highlight and perform a comprehensive study on calibration attacks, a form of adversarial attacks that aim to trap victim models to be heavily miscalibrated without altering their predicted labels, hence endangering the trustworthiness of the models and follow-up decision making based on their confidence. We propose four typical forms of calibration attacks: underconfidence, overconfidence, maximum miscalibration, and random confidence attacks, conducted in both the black-box and white-box setups. We demonstrate that the attacks are highly effective on both convolutional and attention-based models: with a small number of queries, they seriously skew confidence without changing the predictive performance. Given the potential danger, we further investigate the effectiveness of a wide range of adversarial defence and recalibration methods, including our proposed defences specifically designed for calibration attacks to mitigate the harm. From the ECE and KS scores, we observe that there are still significant limitations in handling calibration attacks. To the best of our knowledge, this is the first dedicated study that provides a comprehensive investigation on calibration-focused attacks. We hope this study helps attract more attention to these types of attacks and hence hamper their potential serious damages. To this end, this work also provides detailed analyses to understand the characteristics of the attacks.
@@ 1 Introduction
While recent machine learning models have significantly improved the state-of-the-art performance on a wide range of tasks (Bengio et al., 2021; Vaswani et al., 2017; LeCun et al., 2015; Krizhevsky et al., 2012), these models are often vulnerable and easily deceived by perturbed input (Ren et al., 2020). Adversarial attacks (Ren et al., 2020) have been shown to be a crucial tool to reveal the susceptibility of victim models  (Ibitoye et al., 2019; Zimmermann et al., 2022; Xiao et al., 2023). In the classic setup, adversarial examples are generated by introducing an imperceptible modification to an original datapoint to cause misclassification, where the focus is on trapping victim models to make incorrect predictions.
In this paper, we highlight and provide a comprehensive study on a different type of threats, which we call calibration attacks. The attacks focus on the victim models’ confidence scores without modifying their predicted labels, hence endangering any follow-up decision-making that is based on the victim models’ confidence. We propose to conduct four forms of calibration attacks: underconfidence, overconfidence, maximum miscalibration, and random confidence attacks, which can seriously skew confidence and cause heavy miscalibration, as demonstrated in the reliability diagrams in Figure 1. As we will show and discuss in our study, calibration attacks are insidious and hard to detect. The intrinsic harm is that on the surface the models appear to still make correct decisions, but the level of miscalibration could make the models’ decisions malicious for downstream tasks.
Our specific studies consist of four forms of attacks, span over black-box and white-box setups, attack typical convolutional and attention-based models, and investigate both attack and defence methods. In summary, our main contributions are as follows.
  • To the best of our knowledge, this is the first dedicated study that provides a comprehensive investigation on confidence-focused calibration attacks.
  • We propose to perform four typical forms of calibration attacks and demonstrate their effectiveness and danger to victim models from different perspectives. Detailed insights are provided to understand the characteristics of the attacks and the vulnerability of victim models.
  • We further investigate the effectiveness of a wide range of adversarial defence and recalibration methods, including our proposed defences specifically designed for calibration attacks to mitigate the harm. We hope our work helps attract more attention to these attacks and hence hamper their potential serious damages in applications.
@@
Figure 1: Reliability diagrams of a ResNet-50 classifier (fine-tuned and tested on Caltech-101) before and after the four forms of calibration attacks. Red bars show the average accuracy on the test data binned by confidence scores (15 bins) and the blue bars are the average confidence of samples in each bin. The x-axis represents the bins and y-axis is the accuracy (for red bars) or confidence (for blue bars). The yellow line represents perfect calibration. To have the minimum possible ECE the red bars and blue bars have to completely overlap in each bin (shown in maroon), where no overlap represents miscalibration. Despite the accuracy being unchanged, the miscalibration is severe after the attacks.
@@ 2 Related Work
Among the two typical types of calibration methods, post calibration is applied directly to the predictions of fully trained models at test time, which include classical approaches such as temperature scaling (Guo et al., 2017), Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002), and histogram binning (Zadrozny & Elkan, 2001). Training-based approaches, however, often add bias to help calibrate a model during training (Zhang et al., 2018; Thulasidasan et al., 2019; Kumar et al., 2018; Tomani & Buettner, 2021). In our work we investigate a diverse range of calibration methods against calibration attacks to real the limitations that require them to be overhauled to deal with attacks, including the vulnerability of models on the convolutional architectures (Guo et al., 2017; Minderer et al., 2021) and Transformer-based frameworks (Dosovitskiy et al., 2021). (Refer to Appendix A for a more detailed summary of related work.)
Adversarial Attacks and Training.
Adversarial attacks include black-box and white-box approaches. The former  (Carlini & Wagner, 2017) assume less information about victim models. Many methods are based on gradient estimation through querying the models and finding the finite differences (Bhagoji et al., 2018). White-box attacks often have access to the full details of a victim model such as model architectures and gradients (Goodfellow et al., 2015). Adversarial training and defence approaches have been introduced to improve the robustness of victim models (Stutz et al., 2020; Chen et al., 2022; Qin et al., 2021; Patel et al., 2021; Dhillon et al., 2018). Related to these are works on generating certified robustness guarantees Kumar et al. (2020), where certified radii are generated for the predicted confidence of a smoothed classifier. The research in Emde et al. (2023) further extends this by certifying calibration through generating the worst-case bounds of calibration error, and discuss the importance studying attacks targeting calibration. In this work we propose two defence models against calibration attacks.
Attacking Uncertainty Estimates.
@@
Galil & El-Yaniv (2021) first identified attacks on credible uncertainty estimates as an issue. Their models harm the potential of using uncertainty estimation on a model’s predictions by pushing correct datapoints closer to the decision boundary and incorrect ones further away using confidence without causing incorrect predictions. Zeng et al. (2023) introduce a data poisoning attack designed to alter the training process of models such that a high-confidence region forms around an out-of-domain datapoint, harming the adversary-resistant uncertainty estimates on a set of targeted out-of-domain datapoints while leaving the original labels. Nevertheless, none of the prior works provide a comprehensive study on the calibration attacks, nor do they systematically investigate how well victim models would remain calibrated under such attacks, including under a range of different defence methods. Our work proposes and examines the effects of different attacks, the difficulty of detecting them, and the effectiveness of defence and calibration methods. We provide the study on both convolutional and more recent attention-based models, under both black and white-box approaches. Detailed analyses and insights are additionally discussed.
3 Calibration Attacks
@@
3.1 Objective of Calibration Attacks
Calibration attacks aim to generate adversarial examples to optimize a predefined miscalibration function $\mathcal{M}(\tilde{\mathbf{x}}_{i},k)$ for an adversarial example $\tilde{\mathbf{x}}_{i}$ and the predicted class $k$. As will be detailed below, we propose four forms of calibration attacks where $\mathcal{M}(\tilde{\mathbf{x}}_{i},k)$ takes different implementations. Following the conventional notations, an adversarial example $\tilde{\mathbf{x}}_{i}$ is created by adding noise $\bm{\delta}$ to an input $\mathbf{x}_{i}$: $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\bm{\delta}$, bounded by $\epsilon$ in a $l_{m}$-ball:
@@ 3.1 Objective of Calibration Attacks
$\begin{split}\|\tilde{\mathbf{x}}_{i}-\mathbf{x}_{i}\|_{m}<\epsilon,\>\>m\in\{0,1,\dots,\infty\},\end{split}\vspace{-1mm}$ (1)
where $\epsilon$ controls the amount of allowed perturbation and $m$ corresponds to different norms that may be used. In general, our attacks are based on the most popular view of class-wise calibration (Guo et al., 2017; Kull et al., 2019). For a datapoint $\langle\mathbf{x}_{i},y_{i}\rangle$ in the dataset $\mathcal{D}=\{\langle\mathbf{x}_{n},y_{n}\rangle\}^{N}_{n=1}$, a well calibrated model aims to achieve:
$\begin{split}\mathbb{P}(y_{i}=k\;|\;\hat{p}_{k}(\mathbf{x}_{i})=q_{k})=q_{k},\end{split}\vspace{-1mm}$ (2)
@@
where $q_{k}$ is the confidence of the predicted class $k$ for $\mathbf{x}_{i}$. Any mismatch between the left and right hand sides of the equation creates undesirable miscalibration.
3.2 Four Forms of Calibration Attacks
Building on the above notations and framework, we propose four approaches to cover the most typical variants of calibration attacks.
@@ 3.2 Four Forms of Calibration Attacks
Underconfidence and Overconfidence Attacks (UCA and OCA). These two types of attacks aim to solve the constrained optimization problem involving the miscalibration function $\mathcal{M}(\tilde{\mathbf{x}}_{i},k)$, making a victim model either underconfident or overconfident.
$\mathcal{M}_{\textit{UCA}}{(\tilde{\mathbf{x}}_{i},k)}$ $=\hat{p}_{k}(\tilde{\mathbf{x}}_{i})-\max_{j\neq k}\hat{p}_{j}(\tilde{\mathbf{x}}_{i}),$ (3)
$\mathcal{M}_{\textit{OCA}}(\tilde{\mathbf{x}}_{i},k)$ $=1-\hat{p}_{k}(\tilde{\mathbf{x}}_{i}),$ (4)
$s.t.\>\>\hat{y}(\tilde{\mathbf{x}}_{i})$ $=\hat{y}(\mathbf{x}_{i}).$ (5)
As discussed earlier in introduction, calibration attacks focus on attacking the confidence of victim models but not altering the predicted labels, which is constrained by Eq. 5. By minimizing the loss $\mathcal{M}_{\textit{UCA}}$ or $\mathcal{M}_{\textit{OCA}}$, the attack models maximize calibration errors in these two setups respectively, with the corresponding adversarial examples $\bm{\mathbf{\tilde{X}}}$ :
$\begin{split}\;\mathop{\max}_{\bm{\mathbf{\tilde{X}}}}(\mathbb{P}(y_{i}=k\;|\;\hat{p}_{k}(\tilde{\mathbf{x}}_{i})=q_{k})-q_{k}).\end{split}$ (6)
Algorithm 1 depicts an overview of calibration attacks, including UCA and OCA, but also MMA and RCA that we will introduce below. Unlike the conventional attacks that modify the predicted labels and focus mainly on correctly classified examples, calibration attacks modify confidence and focus on both the originally correctly classified and misclassified instances. As detailed later in Section 4, we will implement calibration attacks in popular black-box  (Andriushchenko et al., 2020) and white-box frameworks (Madry et al., 2018).
1:  Input: A classifier $\mathcal{F}$, input $\mathbf{x}_{i}$, true label $y_{i}$, error bound $\epsilon$, max number of attack iterations $\mathcal{I}$,
attack type $\mathcal{T}$.
2:  Output: Adversarial example $\tilde{\mathbf{x}}_{i}$
[…]
// $k$ is the predicted label for the original (unattacked) input $\mathbf{x}_{i}$.
4:  if $\mathcal{T}$ = UCA or $\mathcal{T}$ = OCA  then
5:     $\mathcal{T^{\prime}}$ = $\mathcal{T}$ ;
6:  else if $\mathcal{T}$ = MMA and $y_{i}\neq k$ then  ;
[…]
7:     $\mathcal{T^{\prime}}$ = OCA
8:  else if $\mathcal{T}$ = MMA and $y_{i}=k$  then
9:     $\mathcal{T^{\prime}}$ = UCA
10:  else if $\mathcal{T}$ = RCA then  ;
[…]
// RCA also considers both UCA and OCA.
11:     $g$ $\xleftarrow{}$ $random(1/K,1.0)$;    ;
// Get a random number in the range $[1/K,1.0]$.
12:     if $g$ > $\hat{p}_{k}(\mathbf{x}_{i})$  then
13:        $\mathcal{T^{\prime}}$ = OCA
14:     else if $g$ < $\hat{p}_{k}(\mathbf{x}_{i})$ then
15:        $\mathcal{T^{\prime}}$ = UCA
16:     end
17:  end
18:  $\tilde{\mathbf{x}}_{i}\leftarrow\mathbf{x}_{i}$; $l_{old}\xleftarrow{}\mathcal{M}_{\mathcal{T^{\prime}}}(\tilde{\mathbf{x}}_{i},k)$
19:  for $i=1$ to $\mathcal{I}$ do
20:     $\bm{\delta}=FindPerturb\;(\tilde{\mathbf{x}}_{i},\epsilon)$ // Find perturbation based on an attack algorithm (e.g., Square Attack) and bound.
21:     $\tilde{\mathbf{x}}_{new}\leftarrow$ $\tilde{\mathbf{x}}_{i}+\bm{\delta}$;
[…]
23:     $l_{new}\leftarrow\mathcal{M}_{\mathcal{T}^{\prime}}(\tilde{\mathbf{x}}_{new},k)$
24:     if ($l_{new}<l_{old}$ and $k_{new}=k$)  then
25:        $\tilde{\mathbf{x}}_{i}\xleftarrow{}\tilde{\mathbf{x}}_{new};l_{old}\xleftarrow{}l_{new}$
26:     end
27:     if ($l_{old}<0.01$) or ($\mathcal{T}$ = RCA and $\hat{p}_{k}(\tilde{\mathbf{x}}_{i})=g$then
28:        Break the for loop
29:     end
30:  end for
Algorithm 1 A Brief Overview of Our Calibration Attack Framework
Maximum Miscalibration Attacks (MMA). We propose MMA with the aim of exploring and understanding more serious scenarios of miscalibration. The main principle of MMA is the aim to perturb (i) all incorrectly classified datapoints to have zero accuracy and 100% confidence, and (ii) all correctly classified attacks to have 100% accuracy but the minimum possible confidence. MMA is a combination of overconfidence and underconfidence attack, as shown in Algorithm 1. Proposition 3.1 below states the property of MMA in terms of the oracle (upper-bound) ECE score that can be achieved in theory, with the proof provided in Appendix B.
Proposition 3.1.
Assume $q$ is the accuracy of a K-way classifier $\mathcal{F}$ on the dataset $\mathcal{D}=\{\langle\mathbf{x}_{n},y_{n}\rangle\}^{N}_{n=1}$. The Maximum Misinformation Attack (MMA) maximizes the expected calibration error (ECE). The upper bound of ECE that can be achieved by MMA is $1-q/K$.